OpenStudy (anonymous):

A triangle has coordinates A (1, 5), B (-2, 1) and C (0, -4). What are the new coordinates if the triangle is dilated with a scale factor of 1/5?

OpenStudy (anonymous):

@ganeshie8 my last question

OpenStudy (anonymous):

@hartnn

ganeshie8 (ganeshie8):

rule for dilation with a scalefactor of \(k\) : \[(x,~y) \longrightarrow (kx~,ky)\]

ganeshie8 (ganeshie8):

simply multiply 1/5 to all the coordinates

ganeshie8 (ganeshie8):

\[(1,5) \longrightarrow(\color{blue}{\frac{1}{5}}*1,~\color{blue}{\frac{1}{5}}*5)~~ =~~ (\frac{1}{5},~1)\]

ganeshie8 (ganeshie8):

see if you can work out remaining two vertices

OpenStudy (anonymous):

1/5 --> (1/5*-2, 1/5 *1) = (-2/5, 1)

OpenStudy (anonymous):

oops i meant (-2,1)

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

ok so this is all i got

OpenStudy (anonymous):

1/5 --> (1/5*1, 1/5*5) = (1/5,1)

OpenStudy (anonymous):

(-2,1) --> (1/5*-2,1/5*1) = (-2/5,1/5)

OpenStudy (anonymous):

(0,-4) --> (1/5*0,1/5*-4) = (0,-4/5)

ganeshie8 (ganeshie8):

Perfect!

OpenStudy (anonymous):

ok am i done here?